# Character-Adapter: Prompt-Guided Region Control for High-Fidelity Character Customization

### v2

@@ 4.4 Ablation study
[…]
 w/o {Prompt-guided segmentation} Region Dynamic 26.9 77.9 47.2
 w/o {Region-level adapter} Image Dynamic 28.4 79.8 53.9
 w/o {Attention dynamic fusion} Region Mask 27.4 78.3 62.2
Character-Adapter (full) Region Dynamic 30.8 86.9 68.3
Region-level adapters. The experimental results in Table 2 show a significant drop in the CLIP-I score when using an image-level adapter. A comparison between Fig. 5 (b) and Fig. 5 (d) demonstrates that the proposed region-level adapters effectively mitigate content fusion issues, thereby improving character consistency.
Attention dynamic fusion. Fig.5 (c) showcases the result without using attention dynamic fusion. The visualization in Fig. 5 (b) reveals a dramatic change in the character’s hairstyle compared to Fig.5 (d). This is because the solid mask only preserves details of certain regions of the reference character and is influenced by the layout image. The result in Table 2 also corroborates the effectiveness of the proposed attention dynamic fusion module.
4.5 User study
We conducted a user study with 20 experts to evaluate Character-Adapter against previous methods. Each expert assessed 50 pairwise comparisons. As shown in Table 3, the results indicate that Character-Adapter achieves a win rate significantly exceeding 50% for both text alignment and character consistency.
Table 3: User study results(%) illustrating the comparison between Character-Adapter and SOTA methods, revealing a notable preference among participants for Character-Adapter in terms of both textual alignment and character consistency.
Methods Models (Character-Adapter vs. *) Text Alignments Character Consistency
@@ 4.5 User study
Finetuning LoRA 92.4 7.6 90.6 9.4
Plug-and-play IP-Adapter 87.5 12.5 70.8 29.2
@@ Appendix A Appendix
We have provided supplementary details regarding our Character-Adapter in this section. Implementation details are illustrated in Section A.1. Limitations and future work are discussed in Sections A.2 and A.3, respectively. Potential societal impacts of our work are examined in Section A.4.
@@ A.1.2 Evaluation metrics
A.2 Limitations and discussion
@@ A.4 Societal impacts
A.5 Additional results
In Fig. 7, we present additional results showcasing the capabilities of Character-Adapter. The proposed framework demonstrates high-fidelity generation of consistent characters and objects across diverse scene contexts. Furthermore, we illustrate that Character-Adapter excels in scenarios involving multiple characters, maintaining high-fidelity of each character throughout the generated images. These results further validate the efficacy and versatility of our proposed approach.

### v4

@@ 4.4 Ablation Study
[…]
 w/o {Prompt-guided segmentation} 26.9 77.9 47.2
 w/o {Region-level adapter} 28.4 79.8 53.9
 w/o {Attention dynamic fusion} 27.4 78.3 62.2
Character-Adapter (full) 30.8 86.9 68.3
@@
Models (Ours vs. *) Text Alignments Character Consistency
@@
LoRA 92.4 7.6 90.6 9.4
IP-Adapter 87.5 12.5 70.8 29.2
@@
Table 3: User study results(%) illustrating the comparison between Character-Adapter and SOTA methods.
Region-level Adapters. The experimental results in Table 2 show a significant drop in the CLIP-I score when removing the region-level adapters. A comparison between Fig. 4 (b) and Fig. 4 (d) demonstrates that the proposed region-level adapters effectively mitigate content fusion issues, thereby improving character consistency. According to Tab. 4, we can observe that as the number of regions increased, our model achieved better performance across all metrics. However, when the number of regions reached 3 and 4, the performance improvements became marginal with no significant growth. Consequently, we opted for 3 regions as the optimal configuration.
Attention Dynamic Fusion. Fig.4 (c) showcases the result without using attention dynamic fusion. The visualization in Fig. 4 (b) reveals a dramatic change in the character’s hairstyle compared to Fig.4 (d). This is because the solid mask only preserves details of certain regions of the reference character and is influenced by the layout image. The result in Tab. 2 also corroborates the effectiveness of the proposed attention dynamic fusion module.
Methods CLIP-T(%) CLIP-I(%) DINO-I(%)
Single Multi Single Multi Single Multi
1-region(IPA) 29.8 26.2 83.6 82.4 59.8 58.0
2-region 29.4 28.3 83.9 82.8 60.2 58.2
3-region(Ours) 30.4 30.2 84.8 84.6 68.1 67.8
4-region 30.2 30.2 84.2 84.8 67.4 66.9
Table 4: Quantitative results (%) of Character-Adapter with different prompt-guided regions, compared with IP-Adapter(IPA). The best results are highlighted in bold.
Method 1-char 2-char 3-char 4-char
Time (s) 7.2 8.78 9.32 10.4
VRAM(MB) 5940 8116 10536 12520
Table 5: Inference time and computational efficiency of Character-Adapter in generating varying numbers of characters (n-char) compared with IP-Adapter(IPA).
4.5 Computational Study
Furthermore, we conduct experiments on the inference time and computing resources of generation 1 to 4 characters with resolution of 768 $\times$ 768. As illustrated in Tab. 5, increasing the number of characters from 1 to 2 results in an average increase in inference time of just 1.07 seconds, which represents a 14.9% increase compared to the inference time for a single character. As the number of characters increases, the computational load and memory usage do not exhibit linear growth. This demonstrates the superior engineering efficiency of Character-Adapter. And the A30 GPU is sufficient for running our model, which demonstrates the superior engineering efficiency of Character-Adapter.
4.6 User Study
We conducted a user study with 20 experts to evaluate Character-Adapter against previous methods. Each expert assessed 50 pairwise comparisons. As shown in Tab. 3, the results indicate that Character-Adapter achieves a win rate significantly exceeding 50% for both text alignment and character consistency.
@@ Appendix A Appendix
We have provided supplementary details regarding our Character-Adapter in this section.
@@ A.1.2 Evaluation metrics
A.2 Segmentation Evaluations
In this subsection, we elucidate the differences between our proposed prompt-guided segmentation and Grounded Segment Anything(Grounded-SAM)[36]. Additionally, we explain our rationale for not selecting existing open-world semantic segmentation models, such as Grounded-SAM.
Segmenting from a pre-trained model exhibits a noticeable disparity between our layout image and the reference image as illustrated in Fig. 6. However, when using Grounded-SAM, the accuracy is so high that we are unable to obtain the entire upper body. Our objective is to ensure minimal overlap between the masks of the three parts we aim to extract, thereby maintaining the integrity and continuity of image features. We conducted an experiment using a reference image for demonstration purposes, which is provided in Fig. 7. When attempting to segment the upper body using Grounded-SAM with the prompt “jacket”, we only capture the jacket region, necessitating the use of a threshold to acquire the complete upper body of the reference image. Conversely, using “upper body” as a prompt results in the model returning the entire body. Prompt-guided segmentation, while providing a rough masked region, enables us to obtain a relatively complete and expected mask. In addition to conducting experiments on SD1.5[37], we also validated the effectiveness of our prompt-guided segmentation approach on Kolors [38]. This confirms that our proposed plug-and-play method is applicable to any text-image cross-attention-based model. As illustrated in the figure, the character’s distinctive features are effectively preserved. This further demonstrates that a well-designed model with strong semantic understanding contributes to generating more precise cross-attention maps, ultimately leading to results with higher consistency. Furthermore, as illustrated in Tab. 6, the use of Grounded-SAM significantly increases GPU memory usage, approximately by 10 GB, leading to a substantial increase in computing resource consumption. Consequently, we employ prompt-guided segmentation to partition characters into multiple distinct regions.
Figure 7: Visual comparison of reference image masks obtaining by our proposed prompt-guided segmentation and Segment Anything
Method Mask IoU(%) CLIP-T(%) CLIP-I(%) DINO-I(%) VRAM(MB) Time(s)
Grounded-SAM 63.5 30.4 84.5 67.9 8549 2.2
Prompt-guided Segmentation 62.5 30.4 84.8 68.1 5592 1.17
Table 6: Quantitative results of Prompt-guided Segmentation and Grounded-SAM.
A.3 Hyperparameter Setup
We performed a comprehensive hyperparameter ablation study using the evaluation datasets described in our paper. By systematically fixing one parameter while adjusting the other, we were able to determine optimal values for both. The results of this analysis are presented in Tab. 7.
$\gamma_{1}$ $\gamma_{2}$ CLIP-T(%$\uparrow$ CLIP-I(%$\uparrow$ DINO-I(%$\uparrow$
0.8 0.7 25.5 79.25 67.4
0.8 0.8 30.4 84.8 68.1
0.8 0.9 29.5 83.4 71.2
0.7 0.8 26.2 74.9 67.87
0.9 0.8 27.8 78.76 68.84
Table 7: Quantitative results (%) of Character-Adapter under different thresholds.
A.4 Limitations and discussion
@@ A.6 Societal impacts
A.7 Additional results
In Fig. 8 and Fig. 9, we present additional results showcasing the capabilities of Character-Adapter. The proposed framework demonstrates high-fidelity generation of consistent characters and objects across diverse scene contexts. Furthermore, we illustrate that Character-Adapter excels in scenarios involving multiple characters, maintaining high-fidelity of each character throughout the generated images. These results further validate the efficacy and versatility of our proposed approach. As shown in Fig. 10, we also demonstrate the effectiveness of our approach across various stylized base models, thereby substantiating the generalizability of our proposed method.
Figure 8: Additional qualitative results of multi characters.
Figure 9: Additional qualitative results of single character.